OpenStudy (anonymous):

need your help.... how to find the intersection of two ellipses given equation 4x^2+9y^2-23=0 and 9x^2+4y^2-36=0.. i'm done getting the all the parameters except the intersection.. help please...

OpenStudy (anonymous):

are you trying to solve these equations?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=solve+ {4+x^2+%2B+9+y^2+-+23+%3D%3D0+%2C+9+x^2+%2B+4+y^2+-+36+%3D+0+}

OpenStudy (anonymous):

\[ \left\{\left\{x\to -2 \sqrt{\frac{58}{65}}, y\to -3 \sqrt{\frac{7}{65}}\right\}, \\ \left\{x\to -2 \sqrt{\frac{58}{65}},y\to 3 \sqrt{\frac{7}{65}}\right\}, \\ \left\{x\to 2 \sqrt{\frac{58}{65}},y\to -3 \sqrt{\frac{7}{65}}\right\}, \\ \left\{x\to 2 \sqrt{\frac{58}{65}},y\to 3 \sqrt{\frac{7}{65}}\right\}\right\} \]